# MultiBLiMP 1.0: A Massively Multilingual Benchmark of Linguistic Minimal Pairs

### v2

@@ 1 Introduction
Large language models (LLMs) are often trained on highly multilingual corpora, which enable users to interact with them in a wide range of languages (Grattafiori et al., 2024; Üstün et al., 2024). Multilingual evaluation of LLMs, however, has mostly focused on their functional linguistic competence through tasks requiring world knowledge and language understanding (Singh et al., 2024a, b), while little work has assessed their formal linguistic competence, i.e. the “knowledge of rules and statistical regularities of a language” (Mahowald et al., 2024). In multilingual contexts, the latter is commonly approximated intrinsically through perplexity (Chang et al., 2024b) or extrinsically through performance in generative tasks such as translation or summarization (Dang et al., 2024). These approaches do not truly disentangle formal from functional competence. Moreover, they are very coarse-grained and do not inform us on which specific constructions a model does (not) master.
@@
To accelerate progress in this direction, we introduce MultiBLiMP 1.0, a massively multilingual benchmark of linguistic minimal pairs covering two types of subject-verb agreement (subject-finite-verb and subject-participle) for number, person, and gender; created automatically using two large-scale linguistic resources: Universal Dependencies Nivre et al. (2016, 2020); de Marneffe et al. (2021) and UniMorph (Batsuren et al., 2022). MultiBLiMP is not only a benchmark, but also a pipeline for the automatic creation of highly multilingual benchmarks (Fig. 2), which can scale to many more linguistic phenomena. Taking subject-finite-verb [^2] Referred to in the following as subject-verb agreement for simplicity. and subject-participle agreement as a use case, we present a first version of the benchmark including more than 128,000 minimal pairs across 101  languages. We use this to evaluate 42 LLMs, finding that linguistic competence is strongly driven by model size and language frequency in training data (Fig. 1), is acquired during pre-training, and can deteriorate as a result of post-training.
@@ 2.1 Targeted Syntactic Evaluation
Minimal pair benchmarks have recently been developed for various other languages, using different methods. Some use existing annotated sentences from UD, some craft templates, and some others are built by translating an English dataset. For Chinese, CLiMP (Xiang et al., 2021) generate their minimal pairs with templates adapted from BLiMP, while SLING (Song et al., 2022) edit a constituency treebank with templates and then verify the results with human annotators. ZhoBLiMP (Liu et al., 2024) employ a mixture of adapting paradigms from English BLiMP, and extracting Chinese-specific paradigms from syntax journals and textbooks. JBliMP (Someya and Oseki, 2023) extract minimal pairs from journal articles on Japanese syntax. BLiMP-NL (Suijkerbuijk et al., 2024) create ten sentence pairs per paradigm by hand for Dutch, as well as 90 more sentences generated by ChatGPT and then hand-checked. LINDSEA (Leong et al., 2023) present minimal pairs for Tamil and Indonesian which are created using prompting with GPT3.5 and GPT-4. Kryvosheieva and Levy (2025) create minimal pairs specifically for Swahili noun class agreement, Hindi split ergativity and Basque verb agreement, each chosen for their complexity. Closer to our approach, RuBLiMP (Taktasheva et al., 2024) create pairs from UD-parsed sentences using perturbation rules. Our approach also leverages UD to identify syntactic relations, which we extend multilingually by setting up a language-agnostic minimal pair creation procedure (§4).
@@
Two multilingual minimal pair benchmarks are known to us. The largest, CLAMS (Mueller et al., 2020), creates pairs for English, German, Hebrew, Italian, and Russian for subject-verb agreement by having native speakers translate the sentences from Marvin and Linzen (2018) where applicable. Gulordava et al. (2018) similarly cover English, Hebrew, Italian, and Russian for various types of agreement using sentences from UD treebanks, in which content words are swapped out with others of the same part-of-speech to test the models’ reliance on semantics. Two other notable related efforts present a multilingual benchmark for linguistic acceptability—for which a model is fine-tuned to predict the grammaticality of a single sentence instead of judging a pair (Warstadt et al., 2019)—ScaLA, which is part of ScandEval (Nielsen, 2023), covering the 5 Scandinavian languages, and MELA (Zhang et al., 2024), covering a diverse of 10 languages.
@@ 3.1 Universal Dependencies
Universal Dependencies Nivre et al. (2016, 2020); de Marneffe et al. (2021) is a multilingual treebank collection, containing 296 treebanks for 168 languages. Besides syntactic dependency relations, UD also contains rich morphological feature annotations that we leverage in our pipeline. An example of a UD-annotated sentence can be found in Fig. 2, step 1. We use UD version 2.15 for the 1.0 release of MultiBLiMP Zeman et al. (2024).
We exclude treebanks of spoken language, historical varieties, as well as certain genres, e.g. technical manuals. Generally, we only exclude treebanks for low-resource languages where absolutely necessary, while for high-resource languages, we also exclude very small treebanks that will be of little additional benefit, but may introduce annotation inconsistencies. A full list of the excluded treebanks with reasons for exclusion can be found in § A.1. We then consider the union of all remaining treebanks for each language.
@@ 4 Pipeline
Our pipeline for creating minimal pairs consists of four stages. First, we extract suitable candidates for a phenomenon using dependency parse trees (§4.1). This allows us to determine the key items of interest for a phenomenon, for example, the subject and verb in subject-verb agreement. Next, we validate whether the agreement phenomenon is present in that language based on UD collocation statistics (§4.2). We then inflect one of the words with respect to a particular feature (e.g. number) to create an agreement violation (§4.3). By replacing the item with this inflected form, we create the ungrammatical counterpart for the minimal pair. Finally, we create a minimal pair dataset for each language, balancing the morphological features and number of pairs (§4.4). We now explain this procedure in more detail; a graphical overview is provided in Figure 2.
@@ 4.2 Agreement Validation
  1. 1.
    $\phi$ must be specified for both elements of the agreement relation, either lexically inherent (la fille est tombée; the girl fell) or morphologically realized (les garçons sont tombés; the boys fell). This way, for example, we rule out subject-verb gender agreement in English, since gender is not specified on verbs. This is easily validated using UniMorph and UD by checking whether both nouns and verbs are annotated for the feature.
  2. 2.
@@
We approach Requirement 2 from a collocational perspective: in a language having strict subject-verb agreement for $\phi$, we expect very high co-occurrence between nouns and verbs of the same feature value, and low co-occurrence between nouns and verbs of contrasting values. This can be expressed as a conditional probability:
@@
Note that it is possible this way for a language to have strict agreement in only one direction. We provide an example of this in Table 1, in which we provide subject-verb number agreement for two languages: Dutch, which has strict number agreement for all conditions, and Turkish, which does not require number agreement for plural subjects. Dutch number agreement is marked as significant for both singular and plural, whereas Turkish agreement only demonstrates agreement for singular items.
@@ 4.3 Inflection
Our inflection procedure has two stages. We start by defining the feature to inflect for, with a candidate word from the filtered treebank of § 4.1. The first stage is to find the corresponding lemma and morphological features of this word in the UD and UM databases, based on its form and the features that were annotated in the UD tree. This may yield multiple candidate lemmas and features, depending on the level of detail of the UD features. In the second stage, we then find all matching rows containing contrasting values of the feature that we inflect for. The inflected forms of these matching rows are the inflection candidates that we use to create the minimal pairs. By considering all opposite values we may create multiple pairs from a single sentence, e.g. He is [..] → He are/am [..].
Number SG PL DU Total
sv vs sv vs sv vs sv vs both
S-Verb $88$ $(74)$ $82$ $(63)$ $73$ $(58)$ $58$ $(43)$ $8$ $(3)$ $3$ $(2)$ $89$ $(79)$ $84$ $(65)$ $90$ $(80)$
S-Participle $35$ $(33)$ $28$ $(21)$ $23$ $(18)$ $19$ $(13)$ $3$ $(0)$ $1$ $(0)$ $35$ $(33)$ $30$ $(22)$ $35$ $(33)$
Gender MASC NEUT FEM Total
sv vs sv vs sv vs sv vs both
S-Verb $31$ $(16)$ $27$ $(16)$ $13$ $(7)$ $13$ $(6)$ $30$ $(13)$ $25$ $(11)$ $34$ $(19)$ $30$ $(16)$ $34$ $(22)$
S-Participle $26$ $(25)$ $25$ $(20)$ $11$ $(8)$ $11$ $(8)$ $20$ $(17)$ $20$ $(12)$ $26$ $(25)$ $26$ $(21)$ $26$ $(25)$
Person 1 2 3 Total
sv vs sv vs sv vs sv vs both
S-Verb $71$ $(42)$ $55$ $(25)$ $63$ $(32)$ $38$ $(15)$ $83$ $(62)$ $69$ $(38)$ $86$ $(66)$ $74$ $(44)$ $87$ $(68)$
S-Participle $2$ $(1)$ 00 $(0)$ $1$ $(1)$ 00 $(0)$ $2$ $(2)$ $1$ $(0)$ $3$ $(2)$ $1$ $(0)$ $3$ $(2)$
Total $100$ $(85)$ $90$ $(66)$ $101$ $(87)$
Table 2: The number of unique languages that yielded at least 10 minimal pairs, for each agreement condition. The number between brackets denotes the number of languages for which the binomial test of §4.2 was significant. SV and VS denote the subject-verb word order.
4.4 Dataset Balancing
To level data imbalances, we limit the number of minimal pairs per agreement condition to 100 items. We obtain these 100 items using a weighted downsampling procedure that also balances sentence-level features. The features we incorporate in this procedure are: subject form, verb form, subject-verb distance, all attractors congruent, and all attractors incongruent. The last two features indicate whether the intervening material between subject and verb contained items that had a congruent or incongruent feature with respect to the agreement feature that we re-inflect (e.g. The keys to the cabinet are). We define the sampling probability $Q$ of a minimal pair as the inverse of the joint probability of its features $P(x)$, assuming feature independence:
$P(x)=\prod_{i}P(x_{i})\hskip 28.45274ptQ(x)=\frac{P(x)^{-1}}{\sum_{i}P({x}_{i})^{-1}}$
The feature probability $P(x_{i})$ is expressed as the relative frequency of the feature. By sampling from $Q$ (without replacements), we can balance all features simultaneously. This results in a dataset that is as balanced as possible for multiple features, ensuring both lexical diversity and an even spread of easier and harder sentences.
5 MultiBLiMP 1.0
We run our minimal pair creation pipeline for two agreement phenomena and three features: subject-verb and subject-participle agreement for number, person, and gender. Within each condition we create sub-conditions based on the inflected feature (e.g. SG $\rightarrow$ PL, 2 $\rightarrow$ 3, etc.), and the order of subject and verb. Our procedure results in 128,321 minimal pairs across 101  languages. The total number of minimal pairs we obtain before balancing is 1.4 million. We provide an aggregated overview of the number of languages per condition in Table 2, and a sample of pairs in § A.2. Subject-verb number agreement is the most common agreement type, with 90 languages, whereas we find subject-participle person agreement for only 3 languages. We present the number of minimal pairs per language in Figure 4. We also provide a detailed breakdown of the language families that are currently covered in MultiBLiMP 1.0 in Figure 3. There is a strong Indo-European (IE) bias present in our minimal pairs, caused both by the over-representation of such languages in UD, as well as our choice to focus on agreement phenomena, which tend to be more present in IE languages. In future iterations of MultiBLiMP, we intend to broaden our coverage by focusing on a more diverse set of phenomena.
Figure 4: Distribution of language families present in MultiBLiMP 1.0. See Appendix B.1 for a detailed version of this figure, including the individual languages.
Subject-Verb Subject-Participle Resources Language Subset
Model Size Version N P G N P G Low Mid High GF Aya EU Eng All #best
Llama3 8B base 84.2 87.8 88.0 89.4 94.0 88.0 77.2 90.3 96.2 89.4 95.2 92.7 99.4 86.9 0
70B base 87.4 91.2 91.1 92.1 97.5 91.2 81.1 93.9 97.8 92.6 97.1 95.5 99.0 90.2 2
70B chat 86.5 90.3 90.3 91.7 97.2 90.6 80.3 93.1 96.9 91.9 96.2 94.6 98.3 89.3 0
Aya 32B chat 82.9 87.8 88.1 87.4 95.1 87.0 75.7 89.4 97.7 89.0 97.3 92.8 98.4 86.4 1
Gemma3 27B base 87.2 91.1 91.3 92.8 96.6 91.7 78.3 96.3 98.0 93.2 97.4 97.1 98.6 90.2 3
27B chat 82.7 87.0 86.3 88.7 95.8 86.7 73.1 92.3 94.4 88.9 93.7 93.3 96.0 85.8 0
OLMo2 32B base 79.8 85.0 80.2 85.7 88.2 80.9 74.4 84.6 92.5 85.1 90.8 87.8 99.5 82.7 0
32B chat 78.2 83.9 79.1 84.2 87.1 80.0 72.5 83.6 91.9 84.0 90.0 86.9 99.1 81.5 0
Qwen3 14B chat 82.2 86.4 86.4 87.8 91.5 86.6 74.1 89.9 94.8 88.2 93.8 92.2 98.3 85.3 0
EuroLLM 9B base 82.7 86.5 87.6 89.1 71.5 89.4 72.6 92.0 95.7 88.9 94.9 96.7 99.4 85.8 0
Goldfish 125M 92.4 95.3 92.2 95.2 98.2 90.9 88.0 95.6 95.9 93.8 95.2 95.8 96.4 93.8 14
@@ 4.4 Dataset Balancing
Table 3: Average accuracies per LLM, split out for different phenomena and language subsets. N stands for Number, P for Person, G for Gender. GF is the language subset of Goldfish languages; Aya the subset of Aya languages; EU the subset of EuroLLM languages; Eng the performance on English. The best performing model per category is denoted in boldface, but we exclude Goldfish models from this ranking as they are only evaluated on a subset of languages. #best denotes the number of languages for which this model was significantly better than the others.
@@ 6.3 Training Corpus Language Distribution
Figure 6: Average accuracy score per language, plotted for Llama3-8B against various other models. The Aya and EuroLLM languages are a subset of the Certain Agreement class, and are highlighted for visualization purposes.
@@ 7.1 General Performance
We report the average accuracy of the largest model for each model family in Table 5. We consider results separately for subject-verb and subject-participle agreement, and then split further by feature between Number, Person, and Gender. We also report results for language subgroups split based on the Common Crawl language frequencies: low-resource for the least frequent 60% languages, mid-resource for the languages in the 60–90% frequency range, and high-resource for the most frequent 10% languages. Finally, we report results for the language subsets of the Goldfish, Aya, and EuroLLM models; performance on English and the overall average. We also denote the number of languages for which a model was significantly better than all other models. [^7] We conduct a McNemar test between the best and second-best model based on the item-level binary grammaticality judgments, with $p<0.1$ as significance threshold.
Llama3-70B-base and Gemma3-27B-base perform best overall, both scoring 90.2% on average. These two models also score best on most of the categories, with the Llama model scoring better on low-resource languages, and Gemma better on mid- and high-resource languages. For all models, performance increases with language frequency, demonstrating much of the variance in linguistic ability can be explained by data disparity. Gemma3-27B-base outperforms the Aya and EuroLLM models on their specific language subsets, validating the claim of Kamath et al. (2025) that its better balance of languages leads to strong multilingual performance. Interestingly, both these models are still being outperformed by the Goldfish model series. The Goldfish models are significantly better on 14 out of 101 languages, whereas Llama3-70B and Gemma3-27B are best for 2 and 3 languages, respectively. We present the best model per language in § B.2, and a breakdown of per-language performance in Appendix B.3.

### v3

@@ 1 Introduction
Large language models (LLMs) are often trained on highly multilingual corpora, which enable users to interact with them in a wide range of languages (Grattafiori et al., 2024; Üstün et al., 2024). Multilingual evaluation of LLMs, however, has mostly focused on their functional linguistic competence through tasks requiring world knowledge and language understanding (Singh et al., 2024, 2025), while little work has assessed their formal linguistic competence, i.e. the “knowledge of rules and statistical regularities of a language” (Mahowald et al., 2024). In multilingual contexts, the latter is commonly approximated intrinsically through perplexity (Chang et al., 2024b) or extrinsically through performance in generative tasks such as translation or summarization (Dang et al., 2024). These approaches do not truly disentangle formal from functional competence. Moreover, they are very coarse-grained and do not inform us on which specific constructions a model does (not) master.
@@
To accelerate progress in this direction, we introduce MultiBLiMP 1.0, a massively multilingual benchmark of linguistic minimal pairs covering two types of subject-verb agreement (subject-finite-verb and subject-participle) for number, person, and gender; created automatically using two large-scale linguistic resources: Universal Dependencies (UD, Nivre et al., 2016, 2020; de Marneffe et al., 2021) and UniMorph (Batsuren et al., 2022). MultiBLiMP is not only a benchmark, but also a pipeline for the automatic creation of highly multilingual benchmarks (Fig.˜2), which can scale to many more linguistic phenomena.
Taking subject-finite-verb [^2] Referred to in the following as subject-verb agreement for simplicity. and subject-participle agreement as a use case, we present a first version of the benchmark including more than 128,000 minimal pairs across 101  languages. We use this to evaluate 42 LLMs, finding that linguistic competence is strongly driven by model size and language frequency in training data (Fig.˜1), is acquired during pre-training, and can deteriorate as a result of post-training.
@@ 2.1 Targeted Syntactic Evaluation
Minimal pair benchmarks have recently been developed for various other languages, using different methods. Some use existing annotated sentences from UD, some craft templates, and some others are built by translating an English dataset.
For Chinese, Xiang et al. (2021, CLiMP) generate minimal pairs with templates adapted from BLiMP, while Song et al. (2022, SLING) edit a constituency treebank with templates and then verify the results with human annotators. Liu et al. (2024, ZhoBLiMP) employ a mixture of adapting paradigms from English BLiMP, and extracting Chinese-specific paradigms from syntax journals and textbooks. Someya and Oseki (2023, JBLiMP) extract minimal pairs from journal articles on Japanese syntax. Suijkerbuijk et al. (2025, BLiMP-NL) create ten sentence pairs per paradigm by hand for Dutch, as well as 90 more sentences generated by ChatGPT and then hand-checked. Leong et al. (2023, LINDSEA) present minimal pairs for Tamil and Indonesian which are created using prompting with GPT-3.5 and GPT-4. Kryvosheieva and Levy (2025) create minimal pairs specifically for Swahili noun class agreement, Hindi split ergativity and Basque verb agreement, each chosen for their complexity.
Closer to our approach, Taktasheva et al. (2024, RuBLiMP) create pairs from UD-parsed sentences using perturbation rules. Our approach too leverages UD to identify syntactic relations, which we extend multilingually by setting up a language-agnostic minimal pair creation procedure (§4).
@@
Two multilingual minimal pair benchmarks are known to us. The largest, CLAMS (Mueller et al., 2020), creates pairs for English, German, Hebrew, Italian, and Russian for subject-verb agreement by having native speakers translate the sentences from Marvin and Linzen (2018) where applicable. Gulordava et al. (2018) similarly cover English, Hebrew, Italian, and Russian for various types of agreement using sentences from UD treebanks, in which content words are swapped out with others of the same part-of-speech to test the models’ reliance on semantics. Two other notable related efforts present a multilingual benchmark for linguistic acceptability—for which a model is fine-tuned to predict the grammaticality of a single sentence instead of judging a pair (Warstadt et al., 2019)—ScaLA, which is part of ScandEval (Nielsen, 2023), covering 5 Scandinavian languages, and MELA (Zhang et al., 2024), covering a diverse set of 10 languages.
@@ 3.1 Universal Dependencies
Universal Dependencies Nivre et al. (2016, 2020); de Marneffe et al. (2021) is a multilingual treebank collection, containing 296 treebanks for 168 languages. Besides syntactic dependency relations, UD also contains rich morphological feature annotations that we leverage in our pipeline. An example of a UD-annotated sentence can be found in Fig.˜2, step 1. We use UD version 2.15 for the 1.0 release of MultiBLiMP Zeman et al. (2024).
We exclude treebanks of spoken language, historical varieties, as well as certain genres, e.g. technical manuals. Generally, we only exclude treebanks for low-resource languages where absolutely necessary, while for high-resource languages, we also exclude very small treebanks that will be of little additional benefit, but may introduce annotation inconsistencies. A full list of the excluded treebanks with reasons for exclusion can be found in §˜A.1. We then consider the union of all remaining treebanks for each language.
@@ 4 Pipeline
Our pipeline for creating minimal pairs consists of four stages. First, we extract suitable candidates for a phenomenon using dependency parse trees (§4.1). This allows us to determine the key items of interest for a phenomenon, like the subject and verb in subject-verb agreement. Next, we validate whether the agreement phenomenon is present in that language based on UD collocation statistics (§4.2). We then inflect one of the words with respect to a particular feature (e.g. number) to create an agreement violation (§4.3). By replacing the item with this inflected form, we create the ungrammatical counterpart for the minimal pair. Finally, we create a minimal pair dataset for each language, balancing the morphological features and number of pairs (§LABEL:sec:balance). We now explain this procedure in more detail; a graphical overview is provided in Figure 2.
@@ 4.2 Agreement Validation
  1. R1:
    $\phi$ must be specified for both elements of the agreement relation, either lexically inherent (la fille est tombée; the girl fell) or morphologically realized (les garçons sont tombés; the boys fell). This way, for example, we rule out subject-verb gender agreement in English, since gender is not specified on verbs. We easily validated this using UniMorph and UD by checking whether both subjects and verbs are annotated for the feature.
  2. R2:
@@
We approach Requirement 2 from a collocational perspective—a similar approach was used by Chaudhary et al. (2020) for grammatical rule detection from treebanks. In a language having strict subject-verb agreement for $\phi$, we expect very high co-occurrence between nouns and verbs of the same feature value, and low co-occurrence between nouns and verbs of contrasting values. This can be expressed as a conditional probability:
@@
$P_{agr}(\phi^{v}|\phi^{n},\text{WO})$
$\phi^{n}$ WO $\phi^{v}$ Dutch Turkish
PL SV PL 0.989 0.161
SG 0.011 0.839
[rgb]0.753,0.753,0.753[rgb]0.0, 0.0, 0.0 PL VS PL 0.990 0.000
SG 0.010 1.000
[rgb]0.753,0.753,0.753[rgb]0.0, 0.0, 0.0 SG SV PL 0.012 0.025
SG 0.988 0.975
[rgb]0.753,0.753,0.753[rgb]0.0, 0.0, 0.0 SG VS PL 0.008 0.018
SG 0.992 0.982
Note that it is possible this way for a language to have strict agreement in only one direction. We show an example in Table 4.2, providing subject-verb number agreement probabilities for two languages: Dutch, which has strict number agreement for all conditions, and Turkish, where plurality is optionally marked on the verb, but singular verbs are commonly used for both singular and plural subjects. Dutch number agreement is marked as significant for both singular and plural, whereas Turkish agreement only demonstrates agreement for singular items.
@@ 4.3 Inflection
Our inflection procedure has two stages. We start by defining the feature to inflect for, with a candidate word from the filtered treebank of §˜4.1. The first stage is to find the corresponding lemma and morphological features of this word in the UD and UM databases, based on its form and the features that were annotated in the UD tree. This may yield multiple candidate lemmas and features, depending on the level of detail of the UD features. In the second stage, we then find all matching rows containing contrasting values of the feature that we inflect for. The inflected forms of these matching rows are the inflection candidates that we use to create the minimal pairs. By considering all opposite values we may create multiple pairs from a single sentence, e.g. He is [..] → He are/am [..].
@@
Table 3: Average accuracies per LLM, split out for different phenomena and language subsets. N stands for Number, P for Person, G for Gender. GF is the language subset of Goldfish languages; Aya the subset of Aya languages; EU the subset of EuroLLM languages; Eng the performance on English. The best performing model per category is denoted in boldface, but we exclude Goldfish models from this ranking as they are only evaluated on a subset of languages. #best denotes the number of languages for which this model was significantly better than the others.7
@@ 6.3 Training Corpus Language Distribution
Figure 6: Average accuracy score per language, plotted for Llama3-8B chat (a) and base (b–d) against various other models. The Aya and EuroLLM languages are a subset of the Certain Agreement class, and are highlighted for visualization purposes.
@@ 7.1 General Performance
We report the average accuracy of the largest model for each model family in Table 4.2. We consider results separately for subject-verb and subject-participle agreement, and then split further by feature between Number, Person, and Gender. We also report results for language subgroups split based on the Common Crawl language frequencies: low-resource for the least frequent 60% languages, mid-resource for the languages in the 60–90% frequency range, and high-resource for the most frequent 10% languages. Finally, we report results for the language subsets of the Goldfish, Aya, and EuroLLM models; performance on English and the overall average. We also denote the number of languages for which a model was significantly better than all other models. [^7] We conduct a McNemar test between the best and second-best model based on the item-level binary grammaticality judgments, with $p<0.1$ as significance threshold.
Llama3-70B-base and Gemma3-27B-base perform best overall, both scoring 90.2% on average. These two models also score best on most of the categories, with the Llama model scoring better on low-resource languages, and Gemma better on mid- and high-resource languages. For all models, performance increases with language frequency, demonstrating much of the variance in linguistic ability can be explained by data disparity. Gemma3-27B-base outperforms the Aya and EuroLLM models on their specific language subsets, validating the claim of Kamath et al. (2025) that its better balance of languages leads to strong multilingual performance. Interestingly, both these models are still being outperformed by the Goldfish model series. The Goldfish models are significantly better on 14 out of 101 languages, whereas Llama3-70B and Gemma3-27B are best for 2 and 3 languages, respectively. We present the best model per language in §˜B.2, and a breakdown of per-language performance in Appendix B.3.